OpenStudy (anonymous):

Solve: 2x + 10 + 26 = 212 Answer 88 98 114 124

OpenStudy (kira_yamato):

x = 88

OpenStudy (anonymous):

2x +36=212 =>2x =176 =>x=88

OpenStudy (ars22):

2x + 10 + 26 = 212 2x + 36 = 212 2x = 212 - 36 2x = 176 x/2 = 176/2 x = 88